# Hybrid Temporal Differential Consistency Autoencoder for Efficient and Sustainable Anomaly Detection in Cyber-Physical Systems

### v1

@@ Abstract
Index Terms:
Theory-Driven Deep Learning, Cyber-Physical Security, Anomaly Detection, Smart Water Distribution Systems, Edge Computing, Environmental Impact
@@ I Introduction
This study makes contributions to three key challenges in anomaly detection for CPS, specifically within the context of water distribution systems. The first challenge is how to leverage the strong time correlations present in the time series data generated by physical processes. The second challenge lies in enhancing machine learning models by integrating physical principles directly into algorithms. The third challenge lies in making machine learning models more efficient and lightweight for edge applications by integrating physical principles and maintaining high performance in anomaly detection while minimizing their computational footprint.
@@ II Related Work
LSTMs (Long Short-Term Memory networks) have proven to be highly effective in capturing temporal dependencies in time series data [11, 12]. Recently, there has been growing interest in applying LSTMs within autoencoders to enhance the detection of anomalies in CPS [13, 14, 15]. While LSTMs are adept at capturing temporal structures, their complexity and resource-intensive nature do not align with our sustainability goals, as they often result in models that are computationally heavy and energy-intensive, and they are not always applicable for edge applications.
Physics-informed neural networks (PINNs) have shown great promise in embedding physical laws directly into neural networks, improving generalization and interpretability in physics-based applications [16, 17, 18]. However, applying PINNs in complex systems like water distribution networks is particularly challenging. These systems are governed by numerous interdependent physical laws, and remodeling the entire system to integrate these laws into the neural network would require significant computational resources and extensive domain expertise, making it impractical for real-world applications [19, 20].
@@
Recent work [1] has proposed a system-theoretic approach inspired by classical embedding theory to analyze the dimensionality of a system’s latent representation. This method introduces physics-inspired consistency principles, which approximate the causal mechanisms governing system dynamics without explicitly enforcing physical laws. The underlying idea is that complex systems in a stable regime exhibit predictable behavior, allowing for well-approximated lower-dimensional embeddings. In contrast, anomalies disrupt these stable relationships, introducing additional complexity into the system.
To incorporate temporal dynamics into the training process, the proposed TDC-AE (Temporal Differential Consistency Autoencoder) employs a Temporal Differential Consistency (TDC) loss. This loss function enforces consistency between the approximated time derivative of static latent variables and the corresponding dynamic latent variables, improving the model’s ability to detect deviations from normal behavior. By integrating lightweight physics-inspired constraints, this method provides an efficient alternative to traditional physics-informed approaches, making it well-suited for real-time anomaly detection in CPS with constrained computational resources.
@@
Finally, sustainability is a critical consideration in designing machine learning models in CPS. Training and deploying these models often demand significant computational resources, leading to higher energy consumption and carbon emissions. As infrastructure systems like water distribution increasingly prioritize sustainability, conventional methods, while effective in terms of detection performance, often overlook energy efficiency. This makes them unsuitable for real-time, large-scale applications where minimizing energy use is essential. This challenge of sustainability was addressed in previous research [21], where edge computing was leveraged to achieve substantial gains in energy efficiency for anomaly detection systems. Building on this, we now focus on refining the internal architecture of the model to further enhance detection performance and maximize energy efficiency.
@@ III Methodology
To address the challenge of distinguishing between deterministic and non-deterministic contributions to system dynamics, we propose a hybrid autoencoder-based approach, hybrid TDC-AE, where the latent space is structured into statistical and deterministic nodes. The deterministic nodes are designed to capture aspects of the system that can be described by physical principles, such as differential equations and conservation laws, while statistical nodes account for residual variability and unmodeled effects. We aim to guide the autoencoder toward a latent representation that aligns with the underlying causal structure of the system.
@@ III-A Dataset
The evaluation is conducted using the publicly available BATADAL dataset, which stands for BATtle of the Attack Detection ALgorithms [22]. This dataset was developed as part of a challenge aimed at comparing the performance of algorithms designed to detect cyber-attacks in water supply systems. It represents the C-Town water distribution system, modeled after a real-world, medium-sized network. The system includes 429 pipes, 388 junctions, 7 storage tanks ($T1-T7$), 11 pumps ($PU1-PU11$) distributed across 5 pumping stations, 5 valves ($V1-V4$), a single reservoir, and nine Programmable Logic Controllers (PLCs). The dataset contains 43 numerical metrics recorded on an hourly basis, with each feature vector representing one hour and consisting of 43 features. These features are readings from the SCADA system, including 7 features representing tank water levels, 12 features related to inlet and outlet pressures for an actuated valve and all pumping stations, and 24 features representing their flow rates and operational status.
@@ III-B Performance Metrics
In order to compare results to literature, some additional performance metrics proposed specifically for the BATADAL challenge [22] were adopted. They read:
  • Performance score $S_{TTD}$, reflecting time-to-detection or the time between a start of an attack and the actual detection of that attack, where $S_{TTD}=1$ being the ideal case in which all attacks are immediately detected, and $S_{TTD}=0$ being the case in which none of the attacks is detected.
  • Classification performance score $S_{CLF}$, defined as the mean value of TPR and TNR.
  • Ranking score $S$, defined as the mean value of the previous two scores – $S_{TTD}$ and $S_{CLF}$.
@@ III-C Data Preprocessing
To ensure our models handle the input data effectively, it is necessary to normalize the features. While standard normalization—subtracting the mean and dividing by the standard deviation—is a typical choice, we employ robust scaling due to its proven effectiveness in previous studies [23, 21]. Robust scaling relies on the median and the interquartile range (IQR), which is the difference between the 75th and 25th percentiles, as expressed in the equation below:
scaled value=original $\text{scaled value}=\frac{\text{original value}-\text{median}}{p_{75}-p_{25}}$ (1)
This approach is more resistant to outliers, preventing extreme values from having an undue influence on the feature distribution. As a result, the scaled features tend to have a standard deviation near 1, without being distorted by the presence of outliers.
@@ IV Temporal Differential Consistency Informed Autoencoders in Water Distribution Networks
As a start we need a way to approximate the first time derivative. The central difference derivative is a numerical method used to approximate the derivative of a function. It is based on evaluating the slope between two points symmetrically spaced around the point of interest. The central difference formula for the first derivative of a function $f(t)$ is:
$\dot{f}(t)\approx\frac{f(t+\Delta t)-f(t-\Delta t)}{2\Delta t}$
where $\Delta t$ is a small step size. This method is preferred over forward or backward difference methods because it provides a higher order of accuracy—specifically, it has an error term that is $O(\Delta t^{2})$, which means the approximation error decreases quadratically as $\Delta t$ decreases [25, 1].
@@ IV-B Integrating Central Difference Calculation into the Latent Space
In alignment with [1], we compute the latent representations at the previous ($t-1$) and next ($t+1$) time steps and apply the central difference method to approximate the time derivative of the static nodes. This derivative is then used to ensure that the dynamic (derivative) nodes in the latent space approximate the central difference derivative of the respective static nodes, considering the time interval $\Delta t$. This approach is shown schematically in Fig. 1. It is important to note that, since each feature is typically min-max scaled, this introduces an abstract time scale, which can potentially affect the interpretability of absolute derivative values.
@@ IV-C Informing Training Process with Temporal Differentials
1:  Initialize: Autoencoder with deterministic latent nodes $\mathbf{z}$, their time derivatives $\dot{\mathbf{z}}$, and statistical latent nodes $\mathbf{s}$, Optimizer, MSE loss function, Weight $\alpha$ TDC loss term.
@@
[…]
@@
[…]
@@
[…]
@@ IV-D Anomaly Estimation based on Autoencoders
Figure 2: Reconstruction error of the BATADAL test dataset on Edge Area 2 with an average window of 7.
Anomaly detection is carried out independently on the test dataset for each edge area using the trained Autoencoders. For each data sample, the reconstruction error is calculated, which forms the basis for detecting anomalies. To smooth out any fluctuations, a moving average filter with a window size of 7 is applied to the reconstruction errors, as suggested by [23]. The threshold is defined as the 95th percentile of the reconstruction errors of the training data. Samples where the reconstruction error surpasses this threshold are classified as anomalies. Fig. 2 provides an example of this approach applied to Edge Area 2, demonstrating how the smoothing filter helps reduce noise and enhance the reliability of anomaly detection.
@@ V-B Detection Performance
No. Reference Approach $S$ $S_{TTD}$ $S_{CLF}$ $F1$ $TPR$ $TNR$ $PPV$ $TP$ $FP$ $TN$ $FN$
@@ VI Discussion & Outlook
In this paper, we introduced hTDC-AE, a novel approach to structuring the latent space of an autoencoder for anomaly detection by incorporating a temporal differential consistency (TDC) loss term during training. Our approach demonstrates stable and reproducible results across multiple training runs. By organizing the latent space into deterministic and statistical nodes, the model shows the potential to bridge the gap between data-driven methods and the integration of physics-inspired principles. Deterministic nodes show indications of encoding features related to physical processes, while statistical nodes capture non-deterministic, abrupt changes such as pump activations. Furthermore, our results suggest that a TDC-based approach enhances the model’s ability to detect not only gradual deviations but also more sudden anomalies, such as tank overflows or abrupt pressure drops, making it applicable to a wider range of real-world scenarios.
Dynamic nodes exhibit strong temporal correlations, approximating the causal link between system states and their derivatives. We attribute the enhanced detection capability to the breakdown of this approximated causal link, which translates into a faster increase in reconstruction error, enabling quicker and more reliable anomaly identification.
Our approach matches the classification performance of state-of-the-art models while significantly improving the time to detect anomalies, a key performance metric in real-world cyber-physical systems, by 3%. This improvement is critical for enhancing system responsiveness in practice. Notably, our model outperforms the leader of the BATADAL challenge without relying on specific domain knowledge, making it a more generalizable solution. Furthermore, our method maintains the same time complexity as conventional autoencoders while reducing the number of fully connected layers from three to two compared to the benchmark. This simplification not only streamlines the architecture but also contributes to a greener, more sustainable model with lower computational requirements.
Our algorithm is highly flexible and can be integrated into any autoencoder for anomaly detection in systems where physical data or deterministic processes are expected. We aim to move beyond conventional statistical fitting of non-linear functions, contributing to the broader challenge of guiding neural networks to learn generalizable and interpretable abstractions that embed meaningful, process-driven relationships.

### v2

@@ I Introduction
This study addresses three key challenges in anomaly detection for CPS, focusing on water distribution systems. The first is leveraging strong time correlations in time series data from physical processes. The second involves enhancing machine learning models by directly integrating physical principles. The third focuses on improving model efficiency for edge applications by maintaining high detection performance while reducing computational demands.
@@ II Related Work
This section follows the problem formulation introduced by [1], which identifies two central challenges in anomaly detection for complex dynamical systems: effective time series modeling and the integration of physical principles into machine learning architectures. LSTMs (Long Short-Term Memory networks) have proven to be highly effective in capturing temporal dependencies in time series data [11, 12]. Recently, there has been growing interest in applying LSTMs within autoencoders to enhance the detection of anomalies in CPS [13, 14, 15]. While LSTMs are adept at capturing temporal structures, their complexity and resource-intensive nature do not align with our sustainability goals, as they often result in models that are computationally heavy and energy-intensive, and they are not always applicable for edge applications. Physics-informed neural networks (PINNs) have shown great promise in embedding physical laws directly into neural networks, improving generalization and interpretability in physics-based applications [16, 17, 18]. However, applying PINNs in complex systems like water distribution networks is particularly challenging. These systems are governed by numerous interdependent physical laws, and remodeling the entire system to integrate these laws into the neural network would require significant computational resources and extensive domain expertise, making it impractical for real-world applications [19, 20].
@@
Recent work [1] has proposed a system-theoretic approach inspired by classical embedding theory to analyze the dimensionality of a system’s latent representation. This method introduces physics-inspired consistency principles, which approximate the causal mechanisms governing system dynamics without explicitly enforcing physical laws. The underlying idea is that complex systems in a stable regime exhibit predictable behavior, allowing for well-approximated lower-dimensional embeddings. In contrast, anomalies disrupt these stable relationships, introducing additional complexity into the system. To incorporate temporal dynamics into the training process, the proposed TDC-AE (Temporal Differential Consistency Autoencoder) employs a Temporal Differential Consistency (TDC) loss. This loss function enforces consistency between the approximated time derivative of static latent variables and the corresponding dynamic latent variables, improving the model’s ability to detect deviations from normal behavior. By integrating lightweight physics-inspired constraints, this method provides an efficient alternative to traditional physics-informed approaches, making it well-suited for real-time anomaly detection in CPS with constrained computational resources.
@@
Sustainability is a key concern in designing machine learning models for CPS. Training and deployment can be resource-intensive, increasing energy use and carbon emissions. As infrastructure systems such as water distribution prioritize sustainability, traditional methods, although effective in detection, often lack energy efficiency and are less suitable for real-time, large-scale applications. This challenge of sustainability was addressed in previous research [21], where edge computing was leveraged to achieve substantial gains in energy efficiency for anomaly detection systems. Building on this, we now focus on refining the internal architecture of the model to further enhance detection performance and maximize energy efficiency.
@@ III Methodology
While TDC aims to enforce meaningful relationships between static and dynamic states in complex dynamical systems, we hypothesize that a more refined representation may be necessary to capture both deterministic and non-deterministic components. Deterministic nodes are organized as static and dynamic pairs, constrained by the TDC loss to model physically meaningful variables. In contrast, statistical nodes are unconstrained and are used to represent residual or discrete behaviors such as pump activation or valve state, which do not follow physical dynamics.
@@ III-A Dataset
The evaluation is conducted using the publicly available BATADAL dataset, which stands for BATtle of the Attack Detection ALgorithms [22]. This dataset was developed as part of a challenge aimed at comparing the performance of algorithms designed to detect cyber-attacks in water supply systems. It represents the C-Town water distribution system, modeled after a real-world, medium-sized network. The system includes 429 pipes, 388 junctions, 7 storage tanks (T1T7T1-T7), 11 pumps (PU1PU11PU1-PU11) distributed across 5 pumping stations, 5 valves (V1V4V1-V4), a single reservoir, and nine Programmable Logic Controllers (PLCs). The dataset contains 43 numerical metrics recorded on an hourly basis, with each feature vector representing one hour and consisting of 43 features. These features are readings from the SCADA system, including 7 features representing tank water levels, 12 features related to inlet and outlet pressures for an actuated valve and all pumping stations, and 24 features representing their flow rates and operational status.
@@ III-B Performance Metrics
In order to compare results to literature, some additional performance metrics proposed specifically for the BATADAL challenge [22] were adopted. The performance score $S_{TTD}$ reflects the time-to-detection, measuring the interval between the start of an attack and its detection, where $S_{TTD}=1$ indicates immediate detection and $S_{TTD}=0$ indicates no detection. The classification performance score $S_{CLF}$ is defined as the mean of the true positive rate (TPR) and true negative rate (TNR). Finally, the overall ranking score $S$ is computed as the mean of $S_{TTD}$ and $S_{CLF}$.
@@ III-C Data Preprocessing
To ensure our models handle the input data effectively, it is necessary to normalize the features. While standard normalization—subtracting the mean and dividing by the standard deviation—is a typical choice, we employ robust scaling due to its proven effectiveness in previous studies [23, 21]. Robust scaling relies on the median and the interquartile range (IQR), which is the difference between the 75th and 25th percentiles. This approach is more resistant to outliers, preventing extreme values from having an undue influence on the feature distribution. As a result, the scaled features tend to have a standard deviation near 1, without being distorted by the presence of outliers.
@@ IV-A Structuring the Latent Space
As a start we need a way to approximate the first time derivative. The central difference derivative is a numerical method used to approximate the derivative of a function. It is based on evaluating the slope between two points symmetrically spaced around the point of interest.
@@ IV-B Integrating Central Difference Calculation into the Latent Space
In alignment with [1], we compute the latent representations at the previous (t1t-1) and next ($t+1$) time steps and apply the central difference method to approximate the time derivative of the static nodes. This derivative is then used to ensure that the dynamic (derivative) nodes in the latent space approximate the central difference derivative of the respective static nodes, considering the time interval $\Delta t$. This approach is shown schematically in Fig. 1. It is important to note that, since each feature is typically min-max scaled, this introduces an abstract time scale, which can potentially affect the interpretability of absolute derivative values.
@@ IV-C Informing Training Process with Temporal Differentials
1:Initialize: Autoencoder with deterministic latent nodes as static/dynamic pairs ($\mathbf{z}$, $\dot{\mathbf{z}}$), and statistical latent nodes $\mathbf{s}$, Optimizer, MSE loss function, Weight $\alpha$ TDC loss term.
@@
11:   $\quad\text{TDC-Loss}\leftarrow\text{MSE}(\Delta_{t}\mathbf{z},\dot{\mathbf{z}_{t}})$
@@
13:   $\quad\text{Rec-Loss}\leftarrow\text{MSE}(\tilde{X}_{t},X_{t})$
@@
15:   $\quad\mathcal{L}=\text{Rec-Loss}+\alpha\cdot\text{TDC-Loss}$
@@ IV-D Anomaly Estimation based on Autoencoders
Anomaly detection is carried out independently on the test dataset for each edge area using the trained Autoencoders. For each data sample, the reconstruction error is calculated, which forms the basis for detecting anomalies. To smooth out any fluctuations, a moving average filter with a window size of 7 is applied to the reconstruction errors, as suggested by [23]. The threshold is defined as the 95th percentile of the reconstruction errors of the training data. Samples where the reconstruction error surpasses this threshold are classified as anomalies.
@@ V-B Detection Performance
No. Reference Approach $S$ $S_{TTD}$ $S_{CLF}$ $F1$ TPRTPR TNRTNR PPVPPV TPTP FPFP TNTN FNFN
@@ VI Discussion & Outlook
In this paper, we introduce hTDC-AE, an extension of TDC-AE [1]. This hybrid approach enhances latent space structuring for anomaly detection by combining a vanilla autoencoder with the previously proposed temporal differential consistency (TDC) loss [1]. Our approach demonstrates stable and reproducible results across multiple training runs. By organizing the latent space into deterministic and statistical nodes, the model shows the potential to bridge the gap between data-driven methods and the integration of physics-inspired principles. Deterministic nodes show indications of encoding features related to physical processes, while statistical nodes capture non-deterministic, abrupt changes such as pump activations. Furthermore, our results suggest that a TDC-based approach enhances the model’s ability to detect not only gradual deviations but also more sudden anomalies, such as tank overflows or abrupt pressure drops, making it applicable to a wider range of real-world scenarios. Dynamic nodes exhibit strong temporal correlations, approximating the causal link between system states and their derivatives. We attribute the enhanced detection capability to the breakdown of this approximated causal link, which translates into a faster increase in reconstruction error, enabling quicker and more reliable anomaly identification. Our approach matches the classification performance of state-of-the-art models while significantly improving the time to detect anomalies, a key performance metric in real-world cyber-physical systems, by 3%. This improvement is critical for enhancing system responsiveness in practice. Notably, our model outperforms the leader of the BATADAL challenge without relying on specific domain knowledge, making it a more generalizable solution. Furthermore, our method maintains the same time complexity as conventional autoencoders while reducing the number of fully connected layers from three to two compared to the benchmark. This simplification not only streamlines the architecture but also contributes to a greener, more sustainable model with lower computational requirements. Our algorithm is highly flexible and can be integrated into any autoencoder for anomaly detection in systems where physical data or deterministic processes are expected. We aim to move beyond conventional statistical fitting of non-linear functions, contributing to the broader challenge of guiding neural networks to learn generalizable and interpretable abstractions that embed meaningful, process-driven relationships.